OpenStudy (anonymous):

The ratio of men to women working for a company is 4 to 5 . If there are 243 employees total, how many women work for the company?

hero (hero):

Ratio of men to total people in the company is \[\frac{4}{9}\] Solve this proportion to find total men working in the company: \[\frac{4}{9} = \frac{x}{243}\]

OpenStudy (anonymous):

so x=180? thanks!

OpenStudy (anonymous):

108*

OpenStudy (anonymous):

then I subtract 108 from 243, correct?

hero (hero):

Smart.

OpenStudy (anonymous):

Okay, great. Thank you :)